A $2,000,000 bond was issued on September 15 2015 with a 5 year maturity. Coupon rate is 5.5% and the present YTM is 5.75% p.From highest to lowest, which financial assets are correctly ordered in terms of their expected return? O a. Ordinary equity,Covenants are designed: O a. to protect the interests of the company. b. to protect debtholders. O c. to protect shareholdersA 10 year bond was issued three years ago. It has a Face Value of $1000 and makes coupon payments of $23 every six months. IfA new company, Jones Ltd, plans to start paying dividends in 4 years time. This first dividend will be $0.90, and thereafter

Hi!

Can you please help me with these questions as my teacher did not provide the answers.

Thanks!
Answer #1
Value of Bond on March15,2018:
Number of semi annual period till maturity 10 (5*2)
Semi annual periods expired till March15,2018 5
Nper Number of semi annual period left 5 (10-5)
Rate Semi annual YTM =(5.75/2)% 2.88%
Pmt Semi annual Coupon payment =(2000000*5.5%)/2 $55,000
Fv Amount to be received on maturity $2,000,000
PV Value of Bond on March15,2018 $1,988,509.75 (Using PV function of excel with Rate=2.88%,Nper=5,Pmt=-55000,Fv=-2000000)
ANSWER:
b.$1,988,509.75
Highest to lowest expected return:
Return and risk are related, higher risk, higher return
Equity has highest risk,Risk of debt is lowest
ANSWER:
a.Ordinary Equity,Preference Shares,Debt
Covenants are designed :
ANSWER:
a.to protect the interest of the company
Face Value $1,000
Yield to maturity 4.60%
Annual Coupon payment =23*2 $46
Coupon payment as percentage of face value=46/1000 4.60%
Since Rate of Coupon and Yield are same,it will sell at par
ANSWER:
a.par
D4=Dividend in year 4=$0.90
g=Dividend growth rate =4%
D5=Dividend in year 5=$0.9*(1+0.04)= $0.936
R=Required rate of return=14%=0.14
P4=Price in year 4 =D5/(R-g)=0.936/(0.14-0.04)= $9.36
Current Market Price =9.36/((1+0.14)^4) $5.54
ANSWER:
c .$5.33
Note, minor difference may be due to approximation
fe =PV(D7,D6,-D8,-D9) D10 Н L. Semi annual periods expired till March15,2018 5 5 (10-5) 6 Nper Number of semi annual period l